# From Prompt to Pipeline: Large Language Models for Scientific Workflow Development in Bioinformatics

### v1

@@ 4.1.1 Workflow Development Using LLMs for Galaxy:
For DeepSeek-V3, we again apply the same instruction-only prompt and are successfully able to generate the steps to develop correct and executable Galaxy workflow for identifying the coding exon with the highest number of SNPs on chromosome 22. Similar to GPT-4o, DeepSeek-V3 recommends the use of the Intersect tool to compute overlaps between SNPs and exon regions. Additionally, the model offers flexibility by suggesting alternative tools for certain tasks. For example, to count the number of SNPs per exon, it proposes using either the Group or Count tool, both valid and appropriate choices. For identifying the exon with the highest SNP count, it recommends the combination of Sort followed by Select, which aligns with standard Galaxy practices. Notably, DeepSeek-V3 also provides supplementary information beyond the core task. It includes guidance on how to export, import, and execute the workflow within the Galaxy platform, along with suggestions for potential public data sources for exons and SNPs. While such details go beyond the minimal workflow requirements, they enhance usability for novice users. Overall, DeepSeek-V3 successfully produces a functionally correct workflow from the instruction-only prompt, demonstrating its capability to support bioinformatics workflow design.
@@
We next evaluate Workflow W2: How to get from peak regions to a list of gene names?. The primary objective is to identify genes that overlap with experimentally derived peak regions and to summarize their distribution across the genome. Following the GTN tutorial, we first upload the dataset GSE37268_mof3.out.hpeak.txt.gz, obtained from the Gene Expression Omnibus (GEO). To annotate the peaks, we retrieve a gene annotation file for the mouse genome from the UCSC Genome Browser and compute overlaps between the peak regions and annotated genes. Finally, we quantify the number of overlapping genes per chromosome to identify regions with enriched regulatory activity. Our analysis shows that chromosome 11 (chr11) contains the highest number of overlapping genes (2164), suggesting notable regulatory significance in this region.
@@
We then proceed with the chain-of-thought (CoT) prompting strategy to further guide GPT-4o in generating a bacterial genome annotation workflow (W5, Table 6). This prompt encourages the model to reason through the task step by step, with the aim of producing a more complete and contextually grounded workflow. While the CoT-generated workflow successfully captures the overarching goals of bacterial genome annotation, namely, structural and functional annotation, quality assessment, and preparation of outputs for visualization, it diverges in several important ways from community-curated standards such as those outlined in the GTN tutorial. Both workflows correctly employ Bakta as the core annotation engine. However, the CoT-generated workflow introduces BUSCO for completeness evaluation, a widely used tool in general genome annotation but not included in the GTN protocol for bacterial genome workflows. More significantly, the GTN tutorial emphasizes a broader approach to structural annotation by incorporating specialized tools such as PlasmidFinder, IntegronFinder, and ISEScan, which detect plasmids, integrons, and insertion sequences, respectively, elements that are essential for accurately characterizing mobile genetic elements in bacterial genomes. The omission of these tools from the generated workflow leads to a structurally incomplete annotation. Additionally, the workflow proposes the use of formatting and indexing tools such as GFF3sort, GFF3 to GenBank, and SAMtools faidx for organizing annotation outputs and enabling visualization. These tools, however, are not available within the Galaxy platform.
In summary, while the generated workflow aligns with the conceptual goals of bacterial genome annotation, it contains key deviations, specifically, the exclusion of critical structural annotation tools and the suggestion of tools not available in Galaxy. As such, the GPT-4o-generated workflow is partially correct but cannot be considered fully usable without further refinement.
@@
The workflow generated by DeepSeek-V3 using an instruction-only prompt demonstrates a generally accurate understanding of the bacterial genome annotation process, yet diverges from the GTN standard in several key areas. For gene prediction, DeepSeek-V3 recommends Prodigal, a well-known tool for prokaryotic CDS identification. For functional annotation, it suggests EggNOG-mapper or BLAST+, both of which are valid in general but require manual output conversion to GFF3, unlike the GTN tutorial’s use of Bakta, which integrates both gene prediction and functional annotation while producing standardized outputs natively compatible with downstream tools like JBrowse. On the other hand, the tools suggested by DeepSeek-V3 for structural annotationPlasmidFinder, IntegronFinder, and ISEScan, are consistent with those used in the GTN tutorial, indicating a correct understanding of the critical elements required to capture plasmid-borne features and mobile genetic elements. However, for non-coding RNA prediction, DeepSeek-V3 proposes the use of Infernal, a powerful RNA homology search tool, which is unfortunately not available on Galaxy, again highlighting a case of tool hallucination. Despite these issues, DeepSeek-V3 correctly outlines the merging of GFF3 outputs and the use of JBrowse for genome visualization, aligning well with the GTN approach in terms of output integration and browser-based annotation exploration.
In comparison to GPT-4o and Gemini 2.5 Flash, DeepSeek-V3 performs moderately well. Like GPT-4o, it hallucinates the availability of tools not supported in Galaxy, reducing practical executability. Gemini, by contrast, stands out for producing a structurally complete and Galaxy-native workflow using only the instruction prompt. While DeepSeek-V3 captures the correct conceptual flow and covers all major analytical components, its output is hindered by limited platform awareness and reliance on tools outside the Galaxy environment. In summary, DeepSeek-V3 offers strong logical grounding but lacks the operational precision demonstrated by Gemini.
@@ 4.1.2 Workflow Development using LLMs for Nextflow:
To evaluate the extent to which LLMs can assist in developing scientific workflows in bioinformatics using Nextflow, we curate a diverse set of representative workflows from the nf-core framework (Table 1). The selected workflows (e.g., fetchngs, methylseq, rnaseq, phyloplace) span multiple domains, including genomics, transcriptomics, epigenomics, and regulatory genomics, thereby enabling a comprehensive and domain-balanced assessment. The nf-core/demo pipeline serves as a foundational example, involving straightforward quality control and adapter trimming of single- or paired-end FASTQ reads, making it useful for testing basic workflow generation capabilities. Fetchngs introduces more advanced functionality by automating the retrieval of sequencing data from public repositories, challenging models to reason about metadata management, remote data access, and parsing of structured sample sheets. Rnaseq and methylseq represent complex, end-to-end pipelines that integrate numerous modules, manage intricate dependencies, and involve advanced data transformations, ideal for evaluating the models’ ability to recommend appropriate toolchains, configurations, and parameters. The inclusion of phyloplace extends the evaluation into the domain of phylogenetic analysis, specifically testing the models’ ability to handle workflows that map query sequences onto a reference phylogenetic tree using specialized inputs and classification tools. These workflows offer a robust and multifaceted benchmark suite to assess how effectively LLMs understand and generate Nextflow-based scientific workflows.
The nf-core/demo (W1 Table 1) workflow is designed to perform a basic bioinformatics data processing task using a minimal dataset. Its core objective is to demonstrate essential steps commonly used in sequencing data preprocessing. First, it conducts quality control of raw FASTQ files using the module FastQC to assess sequence quality metrics. Next, it applies module Seqtk to trim adapter sequences and low-quality bases, improving the quality of the reads. Finally, it aggregates the quality control results using MultiQC, generating a summary report that provides an overview of the data quality before and after trimming. While the workflow is simplified and intended for demonstration, it effectively showcases a typical read preprocessing pipeline involving QC, trimming, and result summarization. We evaluate the workflow using a sample sheet that includes both single-end and paired-end data, available at https://raw.githubusercontent.com/nf-core/test-datasets/viralrecon/samplesheet/samplesheet_test_illumina_amplicon.csv. The workflow is executed using the following command:
nextflow run nf-core/demo -profile docker,test outdir ’demo-results’
@@
We begin by applying the instruction-only prompt (W1, Table 7) to GPT-4o. The model generates a logically structured set of workflow steps, starting with initial quality control using the FastQC module, followed by adapter and quality trimming using tools such as Trim Galore!, fastp, or Trimmomatic. It then suggests performing a second round of quality control to validate the trimming process and concludes with summarizing the results using MultiQC. These steps are largely consistent with those in the official nf-core/demo workflow, which also performs quality control, trimming, and reporting. The primary divergence lies in the trimming step: while the nf-core/demo pipeline uses seqtk, GPT-4o proposes alternative tools that are widely accepted and functionally equivalent within the context of short-read preprocessing. Overall, GPT-4o demonstrates a strong ability to generate a correct and coherent workflow based solely on an instruction-only prompt, indicating a solid understanding of standard practices in NGS data preprocessing with Nextflow.
We then provide the same instruction-only prompt to Gemini 2.5 Flash. The model suggests a detailed and modular workflow design consistent with nf-core best practices. The proposed steps begin with parsing the sample sheet to extract FASTQ file paths using Nextflow channels. It then performs an initial quality assessment using the FastQC module, followed by adapter and quality trimming with either seqtk or Trim Galore!. This is followed by a second round of quality control using FastQC on the trimmed reads, and the final step involves aggregating all quality control outputs into a comprehensive report using MultiQC. While this plan largely mirrors the nf-core/demo workflow, Gemini extends it by emphasizing modularity and offering both seqtk (used in nf-core/demo) and Trim Galore! (which wraps Cutadapt) as valid trimming options. Overall, Gemini 2.5 Flash demonstrates a clear understanding of the workflow structure and appropriately maps tools to each processing stage, showing strong competence in workflow planning from an instruction-only prompt.
@@
We then move to our next workflow nf-core/methylseq, a robust, containerized pipeline designed for comprehensive analysis of bisulfite-converted sequencing (BS-seq) data to facilitate DNA methylation profiling. It encompasses all critical stages of methylation analysis, including data pre-processing, genome alignment, post-alignment processing, and quality control reporting. The workflow automates key tasks such as raw read quality assessment with FastQC, adapter trimming using Trim Galore!, and alignment to a reference genome via either Bismark (with Bowtie2 or HISAT2) or the bwa-meth and MethylDackel toolchain, with optional GPU acceleration supported through Parabricks. In our implementation, we utilize the Bismark-based workflow. The pipeline further includes deduplication of aligned reads, extraction of cytosine methylation metrics, and comprehensive quality evaluations using Qualimap, Preseq, and HS Metrics. Outputs include deduplicated BAM files, detailed methylation calls, and bias reports, all of which are consolidated into a single MultiQC report for streamlined visualization. Additionally, the pipeline generates extensive provenance metadata, including execution logs, software versions, and configuration parameters, ensuring full reproducibility across computing environments. Its modular structure and support for high-performance computing platforms make it well-suited for scalable, large-scale epigenomic studies.
@@
We evaluate the workflow produced by GPT-4o using the role-based prompt. The GPT-4o-generated workflow correctly captures the high-level stages of DNA methylation analysis using bisulfite-converted sequencing (BS-seq) data, beginning with FASTQ quality control and adapter trimming, followed by bisulfite-aware alignment, deduplication, methylation extraction, and MultiQC-based reporting. Importantly, the model successfully identified the correct nf-core module names such as fastqc, trim_galore, bismark/genomeprepare, bismark/align, bismark/deduplicate, and bismark/methylationextractor, indicating that role-based prompting can enhance grounding in actual Nextflow components. However, the generated workflow also exhibits several limitations. First, it does not include configurable options for alternative aligners (e.g., bismark_hisat) or GPU-accelerated execution. Second, while the model included key optional QC modules like qualimap, preseq, and picard/collecthsmetrics, it does not explain when or why to use them, missing opportunities to contextualize their role in specific experimental designs (e.g., capture-based methylation assays). Third, the model omitted support for protocol-specific presets such as RRBS (–rrbs) and PBAT (–pbat), which are natively handled in the nf-core implementation. Finally, although the core logic was structurally correct, some processes like CSV parsing are abstracted as custom steps rather than referencing existing solutions.
To assess the capabilities of Gemini 2.5 Flash in constructing complex scientific workflows, we evaluate its response to an instruction-only prompt (W3, Table 7) for developing a DNA methylation analysis pipeline using bisulfite-converted sequencing (BS-seq) data. The model successfully outlines all critical stages of the workflow, including CSV-based input parsing, raw read quality control, adapter trimming, bisulfite-aware alignment, PCR duplicate removal, methylation extraction, and comprehensive quality control reporting. It recommends appropriate and widely used tools such as FastQC, Trim Galore!, Bismark, MethylDackel, Qualimap, and MultiQC, and its modular structure reflects a good understanding of Nextflow DSL2 practices, including containerization, channel dependencies, and resource management. Furthermore, the workflow is logically organized and aligns well with the stages implemented in the official nf-core/methylseq pipeline. However, Gemini shows some limitations. It does not fully specify the conditional logic or flag handling needed to toggle between these options, nor does it elaborate on the inclusion of protocol-specific parameters. Despite these minor gaps, Gemini 2.5 Flash demonstrates strong proficiency in designing structurally sound and tool-appropriate BS-seq workflows, making it a capable model for assisting in the development of complex, real-world bioinformatics pipelines.
@@
Our next workflow, nf-core/rnaseq, processes raw RNA-seq data into high-quality, analysis-ready outputs. Its main objective is to perform comprehensive quality control, adapter trimming, read alignment, transcript quantification, and reports in a reproducible and scalable manner. The workflow begins with raw FASTQ files and a metadata samplesheet, and takes a reference genome (FASTA) and annotation file (GTF). It includes modules for inferring strand specificity, performing quality checks (e.g., FastQC, MultiQC), trimming adapters (Trim Galore!), filtering contaminants (BBSplit), removing rRNA (SortMeRNA), and aligning reads using STAR or HISAT2. Quantification is performed using Salmon, RSEM, or StringTie, with optional pseudoalignment via Salmon or Kallisto. The pipeline also supports UMI handling, duplicate marking, transcript assembly, and generation of coverage files. Extensive quality metrics are collected using tools such as RSeQC, Qualimap, Preseq, and dupRadar. All results are summarized in an interactive MultiQC report. Designed with portability and reproducibility in mind, the pipeline can be run using Docker, Singularity, or Conda, making it ideal for large-scale or collaborative transcriptomic studies. We execute the workflow using the sample dataset and obtain the desired results. To evaluate the ability of LLMs to generate this complex workflow, we provide an instruction-only prompt (W4, Table 7) and assess whether the models can accurately develop rnaseq pipeline.
We observe that the GPT-4o-generated workflow effectively covers the core analytical stages of RNA-seq data processing, including quality control, adapter trimming, genome alignment, gene/transcript quantification, and differential expression analysis. It accepts a CSV sample sheet with FASTQ links and reference genome inputs, making it flexible and suitable for standard analyses. However, the GPT-4o response has several limitations. The nf-core workflow includes a wider range of features such as pseudoalignment options (e.g., Salmon and Kallisto), UMI handling, strandness autodetection, biotype-level QC, spike-in normalization, and automated GTF/GFF validation, all of which enhance robustness and reproducibility. Moreover, nf-core pipelines benefit from strict input schema validation, containerization, continuous integration testing, and broad community support, making them more reliable for large-scale or production-grade use. The absence of these advanced features in the custom workflow may not affect small or focused studies but could introduce reproducibility challenges and limit flexibility in more complex experimental designs. However, the steps we obtain are fully functional and easygoing.
@@
We conclude our evaluation with the nf-core/phyloplace workflow, which is designed to facilitate efficient and reproducible phylogenetic placement of query sequences onto a reference tree. The workflow operates in two distinct modes: (i) placement-only, where input sequences are directly aligned to a reference multiple sequence alignment and placed onto a fixed phylogenetic tree using likelihood-based methods; and (ii) search-plus-placement, where input sequences from a larger, unfiltered dataset are first screened using profile-based searches to identify candidates for subsequent alignment and placement. After placement, the workflow grafts the query sequences onto the reference tree and performs taxonomic classification, generating annotated phylogenies, placement summaries, and visualizations. Developed using Nextflow DSL2 and fully containerized modules, nf-core/phyloplace ensures portability, scalability, and reproducibility across diverse computing environments. It is particularly well-suited for high-throughput analyses in environmental and metagenomic research, where rapid and accurate phylogenetic integration is critical.
@@
We execute the workflow following the instruction of nf-core/phyloplace. The workflow generates a comprehensive set of output files that summarize the results of the phylogenetic placement and classification process. The workflow produces several key output files organized per sample. These include: (i) placement files containing likelihood-based placements of query sequences on the reference tree; (ii) grafted phylogenetic trees in Newick format with query sequences inserted into the reference topology; (iii) classification tables summarizing the taxonomic or user-defined annotation of each placed sequence; (iv) sequence alignments showing how queries are aligned to the reference multiple sequence alignment; and (v) visualizations and summary reports, such as heat trees and placement statistics, generated via automated reporting tools. A final MultiQC report aggregates log and quality-control metrics across all samples. These outputs facilitate interpretation, visualization, and downstream analysis of the phylogenetic context of newly placed sequences.
@@
To evaluate the ability of LLMs to assist in developing the nf-core/phyloplace workflow, we begin by providing an instruction-only prompt (W5, Table 7) to GPT-4o. The resulting workflow aligns closely with the official nf-core/phyloplace implementation in terms of structure, logic, and tool selection. GPT-4o successfully captures both operational modes, placement-only and search-plus-placement, and accurately identifies the required input components, such as query FASTA files, reference alignments, rooted phylogenetic trees, evolutionary models, HMM profiles, and optional taxonomy files. The model outlines key steps including candidate filtering using hmmsearch, sequence alignment via mafft, phylogenetic placement using epa-ng, tree grafting with gappa, optional taxonomic classification, and summarization with MultiQC. In addition, GPT-4o includes implementation-aware suggestions such as input validation, modular output organization, and execution metadata logging, practices consistent with nf-core standards.

### v2

@@ 4.1.1 Workflow Development Using LLMs for Galaxy:
For DeepSeek-V3, we again apply the same instruction-only prompt and are successfully able to generate the steps to develop correct and executable Galaxy workflow. Similar to GPT-4o, DeepSeek-V3 recommends the use of the Intersect tool to compute overlaps between SNPs and exon regions. Additionally, the model offers flexibility by suggesting alternative tools for certain tasks. For example, to count the number of SNPs per exon, it proposes using either the Group or Count tool, both valid and appropriate choices. For identifying the exon with the highest SNP count, it recommends the combination of Sort followed by Select, which aligns with standard Galaxy practices. Notably, DeepSeek-V3 also provides supplementary information beyond the core task. It includes guidance on how to export, import, and execute the workflow within the Galaxy platform, along with suggestions for potential public data sources for exons and SNPs. While such details go beyond the minimal workflow requirements, they enhance usability for novice users. Overall, DeepSeek-V3 produces a functionally correct workflow from the instruction-only prompt, demonstrating its capability to support bioinformatics workflow design.
@@
We next evaluate Workflow W2: How to get from peak regions to a list of gene names?. The primary objective is to identify genes that overlap with experimentally derived peak regions and to summarize their distribution across the genome. Following the GTN tutorial, we upload the dataset GSE37268_mof3.out.hpeak.txt.gz, obtained from the Gene Expression Omnibus (GEO). To annotate the peaks, we retrieve a gene annotation file for the mouse genome from the UCSC Genome Browser and compute overlaps between the peak regions and annotated genes. Finally, we quantify the number of overlapping genes per chromosome to identify regions with enriched regulatory activity. Our analysis shows that chromosome 11 (chr11) contains the highest number of overlapping genes (2164), suggesting notable regulatory significance in this region.
@@
We then proceed with the CoT prompting strategy to further guide GPT-4o in generating a bacterial genome annotation workflow (W5, Table 6). This prompt encourages the model to reason through the task step by step, with the aim of producing a more complete and contextually grounded workflow. While the CoT-generated workflow successfully captures the overarching goals of bacterial genome annotation, namely, structural and functional annotation, quality assessment, and preparation of outputs for visualization, it diverges in several important ways from community-curated standards such as those outlined in the GTN tutorial. Both workflows correctly employ Bakta as the core annotation engine. However, the CoT-generated workflow introduces BUSCO for completeness evaluation, a widely used tool in general genome annotation but not included in the GTN protocol for bacterial genome workflows. More significantly, the GTN tutorial emphasizes a broader approach to structural annotation by incorporating specialized tools such as PlasmidFinder, IntegronFinder, and ISEScan, which detect plasmids, integrons, and insertion sequences, respectively, elements that are essential for accurately characterizing mobile genetic elements in bacterial genomes. The omission of these tools from the generated workflow leads to a structurally incomplete annotation. Additionally, the workflow proposes the use of formatting and indexing tools such as GFF3sort, GFF3 to GenBank, and SAMtools faidx for organizing annotation outputs and enabling visualization. These tools, however, are not available within the Galaxy platform. In summary, while the generated workflow aligns with the conceptual goals of bacterial genome annotation, it contains key deviations, specifically, the exclusion of critical structural annotation tools and the suggestion of tools not available in Galaxy. As such, the GPT-4o-generated workflow is partially correct but cannot be considered fully usable without further refinement.
@@
The workflow generated by DeepSeek-V3 using the same prompt demonstrates a generally accurate understanding of the bacterial genome annotation process, yet diverges from the GTN standard in several key areas. For gene prediction, DeepSeek-V3 recommends Prodigal, a well-known tool for prokaryotic CDS identification. For functional annotation, it suggests EggNOG-mapper or BLAST+, both of which are valid in general but require manual output conversion to GFF3, unlike the GTN tutorial’s use of Bakta, which integrates both gene prediction and functional annotation while producing standardized outputs natively compatible with downstream tools like JBrowse. On the other hand, the tools suggested by DeepSeek-V3 for structural annotationPlasmidFinder, IntegronFinder, and ISEScan, are consistent with those used in the GTN tutorial, indicating a correct understanding of the critical elements required to capture plasmid-borne features and mobile genetic elements. However, for non-coding RNA prediction, DeepSeek-V3 proposes the use of Infernal, a powerful RNA homology search tool, which is unfortunately not available on Galaxy, again highlighting a case of tool hallucination. Despite these issues, DeepSeek-V3 correctly outlines the merging of GFF3 outputs and the use of JBrowse for genome visualization, aligning well with the GTN approach in terms of output integration and browser-based annotation exploration.
In comparison to GPT-4o and Gemini 2.5 Flash, DeepSeek-V3 performs moderately well. Like GPT-4o, it hallucinates the availability of tools not supported in Galaxy, reducing practical executability. Gemini, by contrast, stands out for producing a structurally complete and Galaxy-native workflow using instruction-only prompt. While DeepSeek-V3 captures the correct conceptual flow and covers all major analytical components, its output is hindered by limited platform awareness and reliance on tools outside the Galaxy environment. In summary, DeepSeek-V3 offers strong logical grounding but lacks the operational precision demonstrated by Gemini.
@@ 4.1.2 Workflow Development using LLMs for Nextflow:
To evaluate the extent to which LLMs can assist in developing scientific workflows in bioinformatics using Nextflow, we curate a diverse set of representative workflows from the nf-core framework (Table 1). The selected workflows span multiple domains, including genomics, transcriptomics, epigenomics, and regulatory genomics, thereby enabling a comprehensive and domain-balanced assessment. The nf-core/demo pipeline serves as a foundational example, involving straightforward quality control and adapter trimming of single- or paired-end FASTQ reads, making it useful for testing basic workflow generation capabilities. Fetchngs introduces more advanced functionality by automating the retrieval of sequencing data from public repositories, challenging models to reason about metadata management, remote data access, and parsing of structured sample sheets. Rnaseq and methylseq represent complex, end-to-end pipelines that integrate numerous modules, manage intricate dependencies, and involve advanced data transformations, ideal for evaluating the models’ ability to recommend appropriate toolchains, configurations, and parameters. The inclusion of phyloplace extends the evaluation into the domain of phylogenetic analysis, specifically testing the models’ ability to handle workflows that map query sequences onto a reference phylogenetic tree using specialized inputs and classification tools. These workflows offer a robust and multifaceted benchmark suite to assess how effectively LLMs understand and generate Nextflow-based scientific workflows.
The nf-core/demo (W1 Table 1) workflow is designed to perform a basic bioinformatics data processing task using a minimal dataset. First, it conducts quality control of raw FASTQ files using the module FastQC to assess sequence quality metrics. Next, it applies module Seqtk to trim adapter sequences and low-quality bases, improving the quality of the reads. Finally, it aggregates the quality control results using MultiQC, generating a summary report that provides an overview of the data quality before and after trimming. We evaluate the workflow using a sample sheet that includes both single-end and paired-end data, available at this link. The workflow is executed using the following command:
nextflow run nf-core/demo -profile docker,test
outdir ’demo-results’
@@
We begin by applying the instruction-only prompt (W1, Table 7) to GPT-4o. The model generates a logically structured set of workflow steps, starting with initial quality control using the FastQC module, followed by adapter and quality trimming using tools such as Trim Galore!, fastp, or Trimmomatic. It then suggests performing a second round of quality control to validate the trimming process and concludes with summarizing the results using MultiQC. These steps are largely consistent with those in the official nf-core/demo workflow. The primary divergence lies in the trimming step: while the nf-core/demo pipeline uses seqtk, GPT-4o proposes alternative tools that are widely accepted and functionally equivalent within the context of short-read preprocessing. Overall, GPT-4o demonstrates a strong ability to generate a correct and coherent workflow based solely on an instruction-only prompt, indicating a solid understanding of standard practices in NGS data preprocessing with Nextflow.
We then provide the same instruction-only prompt to Gemini 2.5 Flash. The model suggests a detailed and modular workflow design. The proposed steps begin with parsing the sample sheet to extract FASTQ file paths using Nextflow channels. It then performs an initial quality assessment using the FastQC module, followed by adapter and quality trimming with either seqtk or Trim Galore!. This is followed by a second round of quality control using FastQC on the trimmed reads, and the final step involves aggregating all quality control outputs into a comprehensive report using MultiQC. While this plan largely mirrors the nf-core/demo workflow, Gemini extends it by emphasizing modularity and offering both seqtk and Trim Galore! (which wraps Cutadapt) as valid trimming options. Overall, Gemini 2.5 Flash demonstrates a clear understanding of the workflow structure and appropriately maps tools to each processing stage, showing strong competence in workflow planning from an instruction-only prompt.
@@
We then move to our next workflow nf-core/methylseq for comprehensive analysis of bisulfite-converted sequencing (BS-seq) data to facilitate DNA methylation profiling. The workflow automates key tasks such as raw read quality assessment with FastQC, adapter trimming using Trim Galore!, and alignment to a reference genome via either Bismark (with Bowtie2 or HISAT2) or the bwa-meth and MethylDackel toolchain, with optional GPU acceleration supported through Parabricks. In our implementation, we utilize the Bismark-based workflow. The pipeline further includes deduplication of aligned reads, extraction of cytosine methylation metrics, and comprehensive quality evaluations using Qualimap, Preseq, and HS Metrics. Outputs include deduplicated BAM files, detailed methylation calls, and bias reports, all of which are consolidated into a single MultiQC report for streamlined visualization. Additionally, the pipeline generates extensive provenance metadata, including execution logs, software versions, and configuration parameters, ensuring full reproducibility across computing environments.
@@
We evaluate the workflow produced by GPT-4o using the role-based prompt. It correctly captures the high-level stages of DNA methylation analysis using bisulfite-converted sequencing (BS-seq) data, beginning with FASTQ quality control and adapter trimming, followed by bisulfite-aware alignment, deduplication, methylation extraction, and MultiQC-based reporting. Importantly, the model successfully identified the correct nf-core module names such as fastqc, trim_galore, bismark/genomeprepare, bismark/align, bismark/deduplicate, and bismark/methylationextractor, indicating that role-based prompting can enhance grounding in actual Nextflow components. However, the generated workflow also exhibits several limitations. First, it does not include configurable options for alternative aligners (e.g., bismark_hisat) or GPU-accelerated execution. Second, while the model included key optional QC modules like qualimap, preseq, and picard/collecthsmetrics, it does not explain when or why to use them, missing opportunities to contextualize their role in specific experimental designs (e.g., capture-based methylation assays). Third, the model omitted support for protocol-specific presets such as RRBS (–rrbs) and PBAT (–pbat), which are natively handled in the nf-core implementation. Finally, although the core logic was structurally correct, some processes like CSV parsing are abstracted as custom steps rather than referencing existing solutions.
To assess the capabilities of Gemini 2.5 Flash in constructing the workflow, we evaluate its response to an instruction-only prompt (W3, Table 7). The model successfully outlines all necessary stages of the workflow, including CSV-based input parsing, raw read quality control, adapter trimming, bisulfite-aware alignment, PCR duplicate removal, methylation extraction, and comprehensive quality control reporting. It recommends appropriate and widely used tools such as FastQC, Trim Galore!, Bismark, MethylDackel, Qualimap, and MultiQC, and its modular structure reflects a good understanding of Nextflow DSL2 practices, including containerization, channel dependencies, and resource management. Furthermore, the workflow is logically organized and aligns well with the stages implemented in the official nf-core/methylseq pipeline. However, Gemini shows some limitations. It does not fully specify the conditional logic or flag handling needed to toggle between these options, nor does it elaborate on the inclusion of protocol-specific parameters. Despite these minor gaps, Gemini 2.5 Flash demonstrates strong proficiency in designing structurally sound and tool-appropriate BS-seq workflows, making it a capable model for assisting in the development of complex, real-world bioinformatics pipelines.
@@
Our next workflow, nf-core/rnaseq, processes raw RNA-seq data into high-quality, analysis-ready outputs. The workflow begins with raw FASTQ files and a metadata samplesheet, and takes a reference genome (FASTA) and annotation file (GTF). It includes modules for inferring strand specificity, performing quality checks (e.g., FastQC, MultiQC), trimming adapters (Trim Galore!), filtering contaminants (BBSplit), removing rRNA (SortMeRNA), and aligning reads using STAR or HISAT2. Quantification is performed using Salmon, RSEM, or StringTie, with optional pseudoalignment via Salmon or Kallisto. The pipeline also supports UMI handling, duplicate marking, transcript assembly, and generation of coverage files. Extensive quality metrics are collected using tools such as RSeQC, Qualimap, Preseq, and dupRadar. All results are summarized in an interactive MultiQC report. To evaluate the ability of LLMs to generate this complex workflow, we provide an instruction-only prompt (W4, Table 7) and assess whether the models can accurately develop rnaseq pipeline.
We observe that the GPT-4o-generated workflow effectively covers the core analytical stages of RNA-seq data processing, including quality control, adapter trimming, genome alignment, gene/transcript quantification, and differential expression analysis. It accepts a CSV sample sheet with FASTQ links and reference genome inputs, making it flexible and suitable for standard analyses. However, the GPT-4o response has several limitations. The nf-core workflow includes a wider range of features such as pseudoalignment options (e.g., Salmon and Kallisto), UMI handling, strandness autodetection, biotype-level QC, spike-in normalization, and automated GTF/GFF validation, all of which enhance robustness and reproducibility. Moreover, nf-core pipelines benefit from strict input schema validation, containerization, continuous integration testing, and broad community support, making them more reliable for large-scale or production-grade use. The absence of these advanced features in the custom workflow may not affect small or focused studies but could introduce reproducibility challenges and limit flexibility in more complex experimental designs. However, the steps we obtain are fully functional.
@@
We conclude our evaluation with the nf-core/phyloplace workflow, which is designed to facilitate efficient and reproducible phylogenetic placement of query sequences onto a reference tree. The workflow operates in two distinct modes: (i) placement-only, where input sequences are directly aligned to a reference multiple sequence alignment and placed onto a fixed phylogenetic tree using likelihood-based methods; and (ii) search-plus-placement, where input sequences from a larger, unfiltered dataset are first screened using profile-based searches to identify candidates for subsequent alignment and placement. After placement, the workflow grafts the query sequences onto the reference tree and performs taxonomic classification, generating annotated phylogenies, placement summaries, and visualizations. Developed using Nextflow DSL2 and fully containerized modules, nf-core/phyloplace ensures portability, scalability, and reproducibility across diverse computing environments. We execute the workflow following the instruction of nf-core/phyloplace. The workflow generates a comprehensive set of output files that summarize the results of the phylogenetic placement and classification process. The workflow produces several key output files organized per sample. These include: (i) placement files containing likelihood-based placements of query sequences on the reference tree; (ii) grafted phylogenetic trees in Newick format with query sequences inserted into the reference topology; (iii) classification tables summarizing the taxonomic or user-defined annotation of each placed sequence; (iv) sequence alignments showing how queries are aligned to the reference multiple sequence alignment; and (v) visualizations and summary reports, such as heat trees and placement statistics, generated via automated reporting tools. A final MultiQC report aggregates log and quality-control metrics across all samples. These outputs facilitate interpretation, visualization, and downstream analysis of the phylogenetic context of newly placed sequences.
@@
To evaluate the ability of LLMs to assist in developing the nf-core/phyloplace workflow, we begin by providing an instruction-only prompt (W5, Table 7) to GPT-4o. The resulting workflow aligns closely with the official nf-core/phyloplace implementation in terms of structure, logic, and tool selection. GPT-4o successfully captures both operational modes, placement-only and search-plus-placement, and accurately identifies the required input components, such as query FASTA files, reference alignments, rooted phylogenetic trees, evolutionary models, HMM profiles, and optional taxonomy files. The model outlines key steps including candidate filtering using hmmsearch, sequence alignment via mafft, phylogenetic placement using epa-ng, tree grafting with gappa, optional taxonomic classification, and summarization with MultiQC. In addition, GPT-4o includes implementation-aware suggestions such as input validation, modular output organization, and execution metadata logging.